# Improved Forward-Forward Contrastive Learning

### v1

@@ 1 Introduction
In the scientific community, there’s a growing interest in finding a biologically plausible alternative to backpropagation. In 2022, Hinton proposed the Forward-Forward (FF) algorithm as one such alternative, drawing inspiration from Boltzmann machines and Noise Contrastive Estimation. Unlike backpropagation, which employs both forward and backward passes, FF relies on two separate forward passes. One pass involves processing positive data, while the other processes negative data. Each layer in the FF model is equipped with its own loss functions. The goal of FF algorithm is to boost the activity within layers for positive data while reducing it for negative data.[Hinton2022].
@@ 2 Proposed Method
In our proposed technique, we have omitted the second and third stages of the original FFCL algorithm. Instead, we utilize two separate instances of the same model, each with the same number and type of layers, including the output classification layer. The weights in the trainable layers are randomly initialized for the two different models. Each set of corresponding trainable layers also has its own loss function, which is utilized for error computation, gradient calculation, and weight update within that layer. Since our method lacks a global scaler representation to monitor the overall performance of the network, we employed the output from one model’s layers as a guiding tool to train the corresponding layer in the second model.
@@ 3 Experiments and Results
The model comprises of four linear layers: an input layer with 784 nodes, two hidden layers with 64 nodes each, and a 10-node output layer. All layers utilize ReLU activations except for the final layer, which employs softmax activation. Cosine embedding serves as the regularization loss for non-classification layers, while binary cross-entropy loss is applied as the classification loss for the last layer. We utilized the Adam optimizer with a learning rate of 0.0001 for weight optimization and trained it for 30 epochs. The MNIST training dataset was leveraged for training purposes, while the test set is used for gathering metrics at the end of each epoch. Positive sampling of data was exclusively carried out, and input images from the dataset are passed through two distinct models concurrently. As the signals propagate forward, the corresponding layers of both models learn simultaneously, obviating the need for backpropagation through layers.

### v3

@@ 1 Introduction
In the scientific community, there’s a growing interest in finding a biologically plausible alternative to backpropagation. In 2022, Hinton proposed the Forward-Forward (FF) algorithm as one such alternative, drawing inspiration from Boltzmann machines and Noise Contrastive Estimation. Unlike backpropagation, which employs both forward and backward passes, FF relies on two separate forward passes. One pass involves processing positive data, while the other processes negative data. Each layer in the FF model is equipped with its own loss functions. The goal of FF algorithm is to boost the activity within layers for positive data while reducing it for negative data [Hinton2022].
@@
Recently Aghagolzadeh and colleagues have introduced a method akin to FFCL, dubbed Forward-Forward with Marginal Supervised Contrastive loss (FFCM). Unlike FFCL’s three-stage training process, FFCM condenses it into two stages. In the first stage, FFCM employs a comparable strategy to train encoder layers as FFCL does. Subsequently, in the second stage, FFCM integrates a classification layer onto the frozen encoder and proceeds with additional training using cross entropy loss for a limited number of epochs [Aghagolzadeh2024]. While FFCM represents an improvement over FFCL, it continues to rely on a two-stage training process. In contrast, our method operates within a single stage, rendering it significantly more efficient than FFCM.
@@ 2 Proposed Method
In our proposed technique, we have omitted the second and third stages of the original FFCL algorithm. Instead, we utilize two separate instances of the same model, each with the same number and type of layers, including the output classification layer. The weights in the trainable layers are randomly initialized for the two different models. Each set of corresponding trainable layers also has a common loss function, which is utilized for error computation, gradient calculation, and weight update within that layer. Since our method lacks a global scaler representation to monitor the overall performance of the network, we employed the output from one model’s layers as a guiding tool to train the corresponding layer in the second model.
@@ 3 Experiments and Results
The model comprises of four linear layers: an input layer with 784 nodes, two hidden layers with 64 nodes each, and a 10-node output layer. All layers utilize ReLU activations except for the final layer, which employs softmax activation. Cosine embedding loss serves as the representation loss for non-classification layers, while binary cross-entropy loss is applied as the classification loss for the last layer. We utilized the Adam optimizer with a learning rate of 0.0001 for weight optimization and trained it for 30 epochs. The MNIST training dataset was leveraged for training purposes, while the test set is used for gathering metrics at the end of each epoch. Positive sampling of data was exclusively carried out, and input images from the dataset are passed through two distinct models concurrently. As the signals propagate forward, the corresponding layers of both models learn simultaneously, obviating the need for backpropagation through layers.